OpenStudy (anonymous):

Fantastic Furniture builds sofas. It costs $2,725 to build 9 sofas and $3,850 to build 14 sofas. Which equation represents the cost, C(x), as a linear function of the number of sofas, x?

OpenStudy (anonymous):

C(x) = 225x - 700 C(x) = 225x + 700 C(x) = 700x - 225 C(x) = 700x + 225

OpenStudy (ganpat):

let a be basic cost and b be manufacturing cost.. so a + 9b = 2725 and a + 14b = 3850 so 14a + (9 * 14)b = 2725 (14) 9a + (9 * 14)b = 3850 (9) subtract 5a = 3500 a = 700 b = 225 so C(x) = 700 + 225x

OpenStudy (anonymous):

thanks!

OpenStudy (ganpat):

w/c..